# Multi-dimensional optimisation of the scanning strategy for the LiteBIRD space mission

### v1

@@ Abstract
Large angular scale surveys in the absence of atmosphere are essential for measuring the primordial $B$-mode power spectrum of the Cosmic Microwave Background (CMB). Since this proposed measurement is about three to four orders of magnitude fainter than the temperature anisotropies of the CMB, in-flight calibration of the instruments and active suppression of systematic effects are crucial. We investigate the effect of changing the parameters of the scanning strategy on the in-flight calibration effectiveness, the suppression of the systematic effects themselves, and the ability to distinguish systematic effects by null-tests. Next-generation missions such as LiteBIRD, modulated by a Half-Wave Plate (HWP), will be able to observe polarisation using a single detector, eliminating the need to combine several detectors to measure polarisation, as done in many previous experiments and hence avoiding the consequent systematic effects. While the HWP is expected to suppress many systematic effects, some of them will remain. We use an analytical approach to comprehensively address the mitigation of these systematic effects and identify the characteristics of scanning strategies that are the most effective for implementing a variety of calibration strategies in the multi-dimensional space of common spacecraft scan parameters. We also present Falcons, a fast spacecraft scanning simulator that we developed to investigate this scanning parameter space.
@@ 1 Introduction
In this paper, we revisit the optimisation of the scanning strategy parameters in a multi-dimensional space with a fast scan simulator that we have developed, and provide clear metrics in order to assess the suppression of systematic effects and calibration for space missions. We also account for a continuously rotating Half-Wave Plate (HWP) that modulates the polarisation of the incoming light. A HWP has been deployed in recent ground-based CMB experiments, such as POLARBEAR and Simons Observatory [12, 13, 14], and is also planned to be equipped on LiteBIRD. By modulating the polarisation with a continuously rotating HWP, it is possible to effectively remove the $1/f$ noise that would otherwise degrade the sensitivity at the large angular scales where the primordial $B$-mode signal becomes prominent. In addition, the pair-differencing between two orthogonal detectors, which is often used for polarisation observations, is no longer necessary with a HWP. Typical systematic effects due to pair-differencing, such as differential gain, differential pointing and differential beam ellipticity, discussed in ref. [15], thus become irrelevant. However, when installing a HWP on a spacecraft, it is necessary to consider other effects, such as whether the polarisation in a region of the sky to which the telescope’s point spread function (i.e. main beam) is pointing would be modulated while it is in transit.
@@ 2.2.2 Constraints on kinetic parameters
Therefore, we assume a situation where the $1/f$ noise with a certain $f_{\rm knee}$ can be sufficiently suppressed by a HWP with revolution frequency $f_{\phi}$, and consider the conditions required for the spacecraft rotation periods. Without a HWP an incoherent polarimetic receiver the polarisation modulation is solely provided by the spacecraft rotation. Such a configuration requires a high spin rate to obtain high modulation efficiency. The specific constraints in this case are discussed in ref. [4], but for a spacecraft with a HWP, the spacecraft’s spin will no longer play a primary role in the modulation. On the other hand, we need to consider whether the HWP can modulate the polarisation in the sky pixel when the telescope’s pointing transits over a sky pixel. The maximum angular velocity at which the pointing sweeps the sky is given by
@@ 2.3 The case of LiteBIRD
In the following analysis, all figures show results that are simulated considering a single boresight detector, i.e., a single detector located at the center of the telescope’s focal plane. [^5] We verified that performing the same study for detectors located at the edge of the FoV in the telescopes, does not change our conclusions regarding the optimisation choice discussed in section 6.
@@ 4.1 Contribution of cross-link factor for systematic effects suppression
4.2 Spin-based map-making
@@
The detailed derivation is written in the LABEL:apd:map-making. In this formalism, if the map of ${{}_{n,m}\tilde{h}}$ is computed in advance under a certain scanning strategy, all calculations in LABEL:eq:map-making_spin are just quadratic operations of the maps, and this is a faster approach than the computationally expensive TOD-based simulations. However, it is difficult for this method to handle time-dependent systematic effects, such as $1/f$ noise or spin-synchronised signals, and TOD simulations are still required in order to correctly assess these.
@@ 4.3 Instance of systematic effects
We introduce a formalism to characterise systematic effects in spin space. This formalism can be used to express some systematic effect without having to consider any time-dependencies.
@@ 5.1 Visibility time of planets
Figure 3 (left) shows the distribution in $\{\alpha,T_{\alpha}\}$ space of the integrated visibility time over the 3 years of the mission duration, which considers all of the observations of the boresight for all of the planets considered. It can be seen that the value of $T_{\alpha}$ has a small impact on the integrated visibility time in the range $T_{\alpha}\lesssim 100$ hours. On the other hand, when $\alpha$ varies, the integrated visibility time reaches its maximum value for $\alpha=\beta=47.5^{\circ}$, and the distribution presents a symmetrical shape centered on this value. This can be explained by the difference in trajectories for each $\alpha$, shown in figure 2. This is because the two parallel lines symmetrically formed across the equator by the shift of the tangent line of the scan due to orbital rotation overlap at the equator when $\alpha=\beta=47.5^{\circ}$. On the other hand, while the planet is traversing the inner gap, that we refer to as the ‘scan pupil’, the planet cannot be observed. The pupil can be seen at the center of the hit-map simulated for a 1-day duration in figure 2. Therefore, scanning strategies with a value of $\alpha$ significantly different from $47.5^{\circ}$ are associated with a lower integrated visibility time.
@@ 5.4 Propagation of cross-link factor to bias
We use the spin-based map-making method presented in section 4 to estimate the measurement bias on $r$, i.e., $\Delta r$ induced by the pointing offset and the HWP non-ideality, and to demonstrate how the value of the cross-link factor can be directly linked to $\Delta r$. A detailed description about the way we estimate $\Delta r$ is given in LABEL:apd:delta_r.
@@
For the pointing offset simulation, we impose the offset parameter $(\rho,\chi)=(1,0)$ arcmin in LABEL:eq:pointing_offset_field and use CMB only map as input. We smooth the CMB map with a symmetric Gaussian beam of $\rm{FWHM}=17.9$ arcmin, simulating the smallest instrumental beam of LiteBIRD, as done in section 2.3. For the simulation of the instrumental polarisation due to the HWP, we impose $\epsilon_{1}=1.0\times 10^{-5}$ and $\phi_{QI}=0$ in LABEL:eq:HWP_IP_field, and use the CMB solar dipole map as input because it is one of the dominant components of temperature as is discussed in ref. [29]. Indeed, CMB and foreground anisotropies can contribute as leakage signals, but CMB anisotropies are sufficiently negligible compared to the dipole. In general, the complexity of the frequency-dependent foregrounds is assessed using specific models and addressed through masking and component separation. This discussion forms a whole topic on itself and we will not consider it here in the optimisation of the scanning strategy. Note that the magnitude of the systematic effect, i.e., $\rho,\chi$ and $\epsilon_{1}$ themselves are not treated here, but only the relative penalty of each scanning strategy is identified. In order to deal with the impact of these systematic effects on the scientific goals, calibration as in ref. [30] and mitigation techniques as in ref. [29] also need to be considered. Notably, the spin-based simulation is about $10^{4}$ times faster than the TOD-based simulation.
@@
Figure 5 (right) shows the distribution of $\Delta r$ due to the instrumental polarisation with the HWP. The distribution has a similar structure to the spin-$(2,0)$ cross-link factor in section 5.3, as expected. This implies that this systematic effect is not suppressed by the reduction of cross-link factors by the HWP rotation, meaning that only the spin-$(2,0)$ cross-link factor acts as a suppression factor. These results indicate that the systematic effect coupled with the spin-$(n,4)$ cross-link factor is suppressed independently of the scanning strategy and that the choice of scanning strategy to reduce spin-$(n,0)$ is important even with the use of HWPs.
@@ 6.2.1 Global survey of the kinetic parameter space
Section 6.2.1 (top panels) shows the $T_{\alpha}$ and $T_{\beta}$ dependence of the spin-$(n,0)$ cross-link factors. Section 6.2.1 (bottom panels) shows the spin-$(2,4)$ cross-link factors calculated for each of the three different telescopes. Note that we zoom into a narrower parameter space region at higher resolution than the upper panels where $T_{\alpha}$ ranges up to 5 hours. For LFT/MFT/HFT, the HWP revolution rates are set to 46/39/61 rpm and $N_{\rm side}$ is set to 128/128/256 because the narrowest FWHM of beams are 23.7/28.0/17.9 arcmin, respectively.  [^13] The size of a HEALPix pixel is $\sqrt{4\pi/N_{\rm pix}}$. When $N_{\rm side}$ is 128/256, the pixel size is 27.9/13.7 arcmin. Parameters other than the spin and precession period are fixed to the values of the standard configuration. Note that the use of the spin rate $\nu_{\beta}$ (instead of $T_{\beta}$), which is often used in spacecraft operations, is given linearly on the horizontal axis of the figures.
@@
For all spin-$(n,0)$ cross-link factors, slowing down $\nu_{\beta}$ does not significantly increase the cross-link factors up to 0.05 rpm, but a 5% increase is observed between 0.04 and 0.05 rpm. On the other hand, the cross-link factor with a HWP contribution increases proportionally to $\nu_{\beta}$, but never monotonically, because it present an oscillating behavior due to the synchronisation of spin and the revolution of the HWP. The value $\nu_{\beta}=0.05$ rpm is the first local minimum that appears when $\nu_{\beta}$ is reduced from $\nu_{\beta}^{\rm upper}$. Therefore, we conclude that $\nu_{\beta}=0.05$ rpm, i.e., $T_{\beta}=20$ min is a compromise point where the spin rate can be slowed down to favor attitude control without dramatically degrading the ability of the scanning strategy to suppress the systematic effects.
We now discuss the optimisation of the precession period. From the viewpoint of gain calibration using the CMB solar dipole, the precession rate as well as the spin rate should be faster to avoid contamination by $1/f$ noise. However, when $\nu_{\beta}$ is 0.05 rpm, the spin-$(n,4)$ cross-link factor is observed to decrease by 15% from $T_{\alpha}^{\rm lower}$ to 3.2 hours of $T_{\alpha}$ in every telescope. On the other hand, the cross-link factor does not change significantly when $T_{\alpha}$ is longer than 3.2 hours (192 minutes), and we conclude that this value is a good trade-off between fast precession period and cross-link factors.
@@ 7.1 Beam reconstruction systematics
Figure 10 (right), displays the amplitude of $\zeta(t)$, defined as the range of variation of the angle made by the direction of scanning in the detector’s frame during one spin cycle. The greater the amplitude, the easier it becomes to solve the degeneracies between unknown systematic effects related to the direction of $\vec{v}_{\rm scan}$. A greater amplitude also facilitate beam-shape reconstruction and characterisation of the sidelobes.
The small value of $T_{\alpha}$ for LiteBIRD is associated with an amplitude of $\zeta$ of approximately $5^{\circ}$, which is a much larger value compared to a mission such as Planck, which had an amplitude of $\zeta$ of 0.25 arcsec. As such, sidelobe reconstruction must be easier with missions like LiteBIRD due to its faster precession. Because of this much larger amplitude of $\zeta(t)$, the presence of unanticipated systematics – such as longer detector time constant or transfer function effects – would be noticeable during the beam reconstruction, leaving some flexibility in order to account for them. However, it is true that the Planck and PICO scanning strategies, with larger $T_{\alpha}$, have the advantage of being able to observe compact sources more frequently over shorter periods.
Figure 10: (left) Illustration of velocity vectors during one spin cycle looking down from the spin axis. Here $x_{\rm FP}$ and $y_{\rm FP}$ represent a focal plane coordinate which indicates the detector reference frame, while $\vec{v}_{\rm prec}$ and $\vec{v}_{\rm spin}$ show the velocity vectors of precession and spin, respectively. The precession motion acts to push the spin axis in the same direction (to the right in the figure) as seen from the spin axis, and the scan velocity vector $\vec{v}_{\rm scan}$, a combination of $\vec{v}_{\rm prec}$ and $\vec{v}_{\rm spin}$, changes the direction across the $x_{\rm FP}$-axis. The angle $\zeta$, referred to as the scanning beam angle, is given by the angle between $\vec{v}_{\rm scan}$ and the $x_{\rm FP}$-axis. (right) Value of the amplitude of $\zeta(t)$ during one spin cycle in the $\{\alpha,T_{\alpha}\}$ space.

### v2

@@ Abstract
Large angular scale surveys in the absence of atmosphere are essential for measuring the primordial $B$-mode power spectrum of the Cosmic Microwave Background (CMB). Since this proposed measurement is about three to four orders of magnitude fainter than the temperature anisotropies of the CMB, in-flight calibration of the instruments and active suppression of systematic effects are crucial. We investigate the effect of changing the parameters of the scanning strategy on the in-flight calibration effectiveness, the suppression of the systematic effects themselves, and the ability to distinguish systematic effects by null-tests. Next-generation missions such as LiteBIRD, modulated by a Half-Wave Plate (HWP), will be able to observe polarisation using a single detector, eliminating the need to combine several detectors to measure polarisation, as done in many previous experiments and hence avoiding the consequent systematic effects. While the HWP is expected to suppress many systematic effects, some of them will remain. We use an analytical approach to comprehensively address the mitigation of these systematic effects and identify the characteristics of scanning strategies that are the most effective for implementing a variety of calibration strategies in the multi-dimensional space of common spacecraft scan parameters. We verify that LiteBIRD’s standard configuration yields good performance on the metrics we studied. We also present Falcons.jl, a fast spacecraft scanning simulator that we developed to investigate this scanning parameter space.
@@ 1 Introduction
In this paper, we revisit the optimisation of the scanning strategy parameters in a multi-dimensional space with a fast scan simulator that we have developed, and provide clear metrics in order to assess the suppression of systematic effects and calibration for space missions. We also account for a continuously rotating Half-Wave Plate (HWP) that modulates the polarisation of the incoming light. A HWP has been deployed in recent ground-based CMB experiments, such as POLARBEAR and Simons Observatory [12, 13, 14], and is also planned to be equipped on LiteBIRD. By modulating the polarisation with a continuously rotating HWP, it is possible to effectively remove the $1/f$ noise that would otherwise degrade the sensitivity at the large angular scales where the primordial $B$-mode signal becomes prominent. In addition, the pair-differencing between two orthogonal detectors, which is often used for polarisation observations, is no longer necessary with a HWP. Typical systematic effects due to pair-differencing, such as differential gain, differential pointing and differential beam ellipticity, discussed in ref. [15], thus become irrelevant. However, when installing a HWP on a spacecraft, it is necessary to consider other effects, such as whether the polarisation in a region of the sky to which the telescope’s point spread function (i.e. main beam) is pointing would be modulated while it is in transit, that is, whether the HWP rotation is sufficiently faster than the sky scanning motion.
@@ 2.2.2 Constraints on kinetic parameters
Therefore, we assume a situation where the $1/f$ noise with a certain $f_{\rm knee}$ can be sufficiently suppressed by a HWP with revolution frequency $f_{\phi}$, and consider the conditions required for the spacecraft rotation periods. Without a HWP an incoherent polarimetic receiver the polarisation modulation is solely provided by the spacecraft rotation. Such a configuration requires a high spin rate to obtain high modulation efficiency. The specific constraints in this case are discussed in ref. [4], but for a spacecraft with a HWP, the spacecraft’s spin will no longer play a primary role in the modulation. On the other hand, we need to consider whether the HWP can modulate the polarisation in the sky pixel when the telescope’s pointing transits over a sky pixel, i.e., whether the HWP rotation is fast enough for the sky scan motion. The maximum angular velocity at which the pointing sweeps the sky is given by
@@ 2.3 The case of LiteBIRD
In the following analysis, all figures show results that are simulated considering a single boresight detector, i.e., a single detector located at the center of the telescope’s focal plane. We verified that performing the same study for detectors located at the edge of the FoV in the telescopes, does not change our conclusions regarding the optimisation choice discussed in section 6. A detailed discussion of the detectors non located at the boresight can be found in section B.2.
@@ 4.1 Contribution of cross-link factor for systematic effects suppression
4.2 Map-based simulation using the spin formalism
@@
A detailed derivation is given in LABEL:apd:map-making. This formalism has applications beyond the one presented in ref. [17] and can be generalised to allow for a HWP to be taken into account. We verify explicitly in LABEL:apd:TOD_comparison that this new method gives results which are consistent with TOD-based binning map-making simulations. Also, here the demodulation is done in the vector part of the right side of LABEL:eq:map-making_spin as we can see in LABEL:eq:map-making_TOD which is describing TOD-based binning map-maker. This method is different from the method used to obtain independent TOD of $I$, $Q$ and $U$ with the lock-in technique using HWP, and its implementation into the map-based simulation is left for future study. [^7] It has not yet been determined what demodulation method will be used in the case of LiteBIRD. However, the time-independent systematic effects discussed in the next section can be accounted for by cross-link factors, so the conclusions of this paper based on them are not dependent on the demodulation method used in map-making.
In this formalism, if the map of ${{}_{n,m}\tilde{h}}$ is computed in advance under a certain scanning strategy, all calculations in LABEL:eq:map-making_spin are just quadratic operations of the maps, and this is a faster approach than the computationally expensive TOD-based simulations. However, it is difficult for this method to handle time-dependent systematic effects, such as $1/f$ noise or scan-synchronised signals, and TOD simulations are still required in order to correctly assess these.
@@ 4.3 Instance of systematic effects
We introduce a formalism to characterise systematic effects in spin space. This formalism can be used to express some systematic effects without having to consider any time-dependencies.
@@ 5.1 Visibility time of planets
Figure 3 (left) shows the distribution in $\{\alpha,T_{\alpha}\}$ space of the integrated visibility time over the 3 years of the mission duration, which considers all of the observations of the boresight for all of the planets considered. There is no significant difference in the shape of the distribution, including the position of the peaks, among the planets considered. The peak heights are as follows: Mars, 0.77 hours; Jupiter, 1.00 hour; Saturn, 0.87 hours; and Neptune, 0.83 hours. It can be seen that the value of $T_{\alpha}$ has a small impact on the integrated visibility time in the range $T_{\alpha}\lesssim 100$ hours. On the other hand, when $\alpha$ varies, the integrated visibility time reaches its maximum value for $\alpha=\beta=47.5^{\circ}$, and the distribution presents a symmetrical shape centered on this value. This can be explained by the difference in trajectories for each $\alpha$, shown in figure 2. This is because the two parallel lines symmetrically formed across the equator by the shift of the tangent line of the scan due to orbital rotation overlap at the equator when $\alpha=\beta=47.5^{\circ}$. On the other hand, while the planet is traversing the inner gap, that we refer to as the ‘scan pupil’, the planet cannot be observed. The pupil can be seen at the center of the hit-map simulated for a 1-day duration in figure 2. Therefore, scanning strategies with a value of $\alpha$ significantly different from $47.5^{\circ}$ are associated with a lower integrated visibility time.
@@ 5.4 Propagation of cross-link factor to bias
We use the map-based method using the spin formalism presented in section 4 to estimate the measurement bias on $r$, i.e., $\Delta r$ induced by the pointing offset and the HWP non-ideality, and to demonstrate how the value of the cross-link factor can be directly linked to $\Delta r$. A detailed description about the way we estimate $\Delta r$ is given in LABEL:apd:delta_r.
@@
For the pointing offset simulation, we impose the offset parameter $(\rho,\chi)=(1,0)$ arcmin in LABEL:eq:pointing_offset_field and use CMB only map as input. We smooth the CMB map with a symmetric Gaussian beam of $\rm{FWHM}=17.9$ arcmin, simulating the smallest instrumental beam of LiteBIRD, as done in section 2.3. For the simulation of the instrumental polarisation due to the HWP, we impose $\epsilon_{1}=1.0\times 10^{-5}$ and $\phi_{QI}=0$ in LABEL:eq:HWP_IP_field, and use the CMB solar dipole map as input because it is one of the dominant components of temperature as is discussed in ref. [29]. Indeed, CMB and foreground anisotropies can contribute as leakage signals, but CMB anisotropies are sufficiently negligible compared to the dipole. In general, the complexity of the frequency-dependent foregrounds is assessed using specific models and addressed through masking and component separation. This discussion forms a whole topic on itself and we will not consider it here in the optimisation of the scanning strategy. Note that the magnitude of the systematic effect, i.e., $\rho,\chi$ and $\epsilon_{1}$ themselves are not treated here, but only the relative penalty of each scanning strategy is identified. In order to deal with the impact of these systematic effects on the scientific goals, calibration as in ref. [30] and mitigation techniques as in ref. [29] also need to be considered. Notably, the map-based simulations using the spin formalism shown in sections 4.2 and LABEL:apd:spin_space_analysis is about $10^{4}$ times faster than the TOD-based simulation.
@@
Figure 5 (right) shows the distribution of $\Delta r$ due to the instrumental polarisation with the HWP. The distribution has a similar structure to the spin-$(2,0)$ cross-link factor in section 5.3, as expected. This implies that this systematic effect is not suppressed by the reduction of cross-link factors by the HWP rotation, meaning that only the spin-$(2,0)$ cross-link factor acts as a suppression factor. The justification of the cross-link factor dependency of the HWP non-ideality can be found in LABEL:apd:HWP_sys. These results indicate that the systematic effect coupled with the spin-$(n,4)$ cross-link factor is suppressed independently of the scanning strategy and that the choice of scanning strategy to reduce spin-$(n,0)$ is important even with the use of HWPs.
@@ 6.2.1 Global survey of the kinetic parameter space
Section 6.2.1 (top panels) shows the $T_{\alpha}$ and $T_{\beta}$ dependence of the spin-$(n,0)$ cross-link factors. Section 6.2.1 (bottom panels) shows the spin-$(2,4)$ cross-link factors calculated for each of the three different telescopes. Note that we zoom into a narrower parameter space region at higher resolution than the upper panels where $T_{\alpha}$ ranges up to 5 hours. For LFT/MFT/HFT, the HWP revolution rates are set to 46/39/61 rpm and $N_{\rm side}$ is set to 128/128/256 because the narrowest FWHM of beams are 23.7/28.0/17.9 arcmin, respectively. [^13] The size of a HEALPix pixel is $\sqrt{4\pi/N_{\rm pix}}$. When $N_{\rm side}$ is 128/256, the pixel size is 27.5/13.7 arcmin. Parameters other than the spin and precession period are fixed to the values of the standard configuration. Note that the use of the spin rate $\nu_{\beta}$ (instead of $T_{\beta}$), which is often used in spacecraft operations, is given linearly on the horizontal axis of the figures.
@@
For all spin-$(n,0)$ cross-link factors, slowing down $\nu_{\beta}$ does not significantly increase the cross-link factors up to 0.05 rpm, but a 5% increase is observed between 0.04 and 0.05 rpm. On the other hand, the cross-link factor with a HWP contribution increases proportionally to $\nu_{\beta}$, but never monotonically, because it presents an oscillating behavior due to the synchronisation of spin and the revolution of the HWP. The value $\nu_{\beta}=0.05$ rpm is the first local minimum that appears when $\nu_{\beta}$ is reduced from $\nu_{\beta}^{\rm upper}$. Therefore, we conclude that $\nu_{\beta}=0.05$ rpm, i.e., $T_{\beta}=20$ min is a compromise point where the spin rate can be slowed down to favor attitude control without dramatically degrading the ability of the scanning strategy to suppress the systematic effects.
We now discuss the optimisation of the precession period. From the viewpoint of gain calibration using the CMB solar dipole, the precession rate as well as the spin rate should be faster to avoid contamination by $1/f$ noise. However, when $\nu_{\beta}$ is 0.05 rpm, the spin-$(n,4)$ cross-link factor is observed to decrease by 15% from $T_{\alpha}^{\rm lower}$ to $T_{\alpha}=3.2$ hours in every telescope. On the other hand, the cross-link factor does not change significantly when $T_{\alpha}$ is longer than 3.2 hours (192 minutes), and we conclude that this value is a good trade-off between fast precession period and cross-link factors.
@@ 7.1 Beam reconstruction systematics
Figure 10 (right), displays the total amplitude of $\zeta(t)$, defined as the range of variation of the angle made by the direction of scanning in the detector’s frame during one spin cycle. The greater the amplitude, the easier it becomes to solve the degeneracies between unknown systematic effects related to the direction of $\vec{v}_{\rm scan}$. A greater amplitude also facilitate beam-shape reconstruction and characterisation of the sidelobes.
The small value of $T_{\alpha}$ for LiteBIRD is associated with an total amplitude of $\zeta$ of approximately $5^{\circ}$, which is a much larger value compared to a mission such as Planck, which had a total amplitude of $\zeta$ of 0.5 arcsec. As such, sidelobe reconstruction must be easier with missions like LiteBIRD due to its faster precession. Because of this much larger total amplitude of $\zeta(t)$, the presence of unanticipated systematics – such as longer detector time constant or transfer function effects – would be noticeable during the beam reconstruction, leaving some flexibility in order to account for them. However, it is true that the Planck and PICO scanning strategies, with larger $T_{\alpha}$, have the advantage of being able to observe compact sources more frequently over shorter periods.
Figure 10: (left) Illustration of velocity vectors during one spin cycle looking down from the spin axis. Here $x_{\rm FP}$ and $y_{\rm FP}$ represent a focal plane coordinate which indicates the detector reference frame, while $\vec{v}_{\rm prec}$ and $\vec{v}_{\rm spin}$ show the velocity vectors of precession and spin, respectively. The precession motion acts to push the spin axis in the same direction (to the right in the figure) as seen from the spin axis, and the scan velocity vector $\vec{v}_{\rm scan}$, a combination of $\vec{v}_{\rm prec}$ and $\vec{v}_{\rm spin}$, changes the direction across the $x_{\rm FP}$-axis. The angle $\zeta$, referred to as the scanning beam angle, is given by the angle between $\vec{v}_{\rm scan}$ and the $x_{\rm FP}$-axis. (right) Value of the total amplitude of $\zeta(t)$ during one spin cycle in the $\{\alpha,T_{\alpha}\}$ space.